your sock draw has 18 folded pair of socks, 8 pair of white socks, 6 pair of black socks, and 4 pair of blue socks. what is the probability, without looking that you will select and remove a black pair, then select either a white or blue pair?
Answer by solve_for_x(190)

The probability of selecting a black pair first is:

P(black) = (number of black pairs) / (total number of pairs)

P(black) = 6/18 = 1/3

The, with only 17 pairs remaining, and 8 + 4 = 12 pairs that are either
white or blue, the probability of selecting either a white or a blue
pair is:

P(white or blue, after black) = (number of white or blue pairs) / (number of pairs remaining)

P(white or blue, after black) = 12/17

The probability of both events happening in succession is:

P(black, followed by white or blue) = (1/3)(12/17) = 4/17